OpenStudy (anonymous):

what is the slope of the lines through the points (-2,4) (-1,-1)

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) in this case y2=-1 y1=4 x2=-1 x1=-2 does that help?

OpenStudy (anonymous):

you need to use the slope formula: \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (skullpatrol):